Given two vector vector a 4.0i^-3.0j^ and vector b=6.0i^+8.0j^? find the magnotude and direction (with +x axis )of (a)vector a(b)vector b(c) vector b-a(d)vector a-b.
 
 

Grade:12th pass

1 Answers

Saurabh Koranglekar
5 years ago
magnitude –> (x2 + y2)0.5
and direction → Tan-1(y/x)

a.​ magnitude=(42+ (-3)2)0.5= 5; direction= tan-1(-3/4)
b.magnitude=(62+ (8)2)0.5= 10; direction= tan-1(8/6)
c. magnitude=​(22+ (11)2)0.5= (125)0.5; direction= tan-1(11/2)
d. magnitude=((-2)2+ (-11)2)0.5= (125)0.5; direction= tan-1(11/2)